Solve: |2x +3| -6 = 11
Answer by lwsshak3(11628)

Solve: |2x +3| -6 = 11
**
|2x +3| -6 = 11
|2x+3|=11+6=17
for (2x+3)>0
2x+3=17
2x=14
x=7
..
for (2x+3)<0
-|2x+3|=17
-2x-3=17
-2x=20
x=-10